# A Comprehensive Survey of Speculative Decoding

### v2

@@ C.1 Details of Spec-Bench
Moreover, Spec-Bench incorporates three further subtasks – translation, question answering, and mathematical reasoning – to provide a thorough evaluation of Speculative Decoding’s speedup capabilities in diverse contexts. We utilized WMT14 DE-EN, Natural Questions, and GSM8K Cobbe et al. (2021) as the primary datasets for these tasks, respectively. We randomly selected 80 instances from each subtask’s test set for evaluation.
C.2 Implement Details
@@
$T=0$ Autoregressive Decoding 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 36.74±${}_{\pm 0.31}$ 1.00$\times$
Lookahead Fu et al. (2024) 1.15$\times$±${}_{\pm 0.01}$ 0.98$\times$±${}_{\pm 0.02}$ 1.07$\times$±${}_{\pm 0.02}$ 1.06$\times$±${}_{\pm 0.02}$ 1.32$\times$±${}_{\pm 0.02}$ 1.03$\times$±${}_{\pm 0.02}$ 40.64±${}_{\pm 0.26}$ 1.11$\times$
REST He et al. (2023) 1.49$\times$±${}_{\pm 0.02}$ 1.23$\times$±${}_{\pm 0.04}$ 1.26$\times$±${}_{\pm 0.03}$ 1.39$\times$±${}_{\pm 0.04}$ 1.34$\times$±${}_{\pm 0.03}$ 1.71$\times$±${}_{\pm 0.05}$ 51.12±${}_{\pm 0.78}$ 1.39$\times$
PLD Saxena (2023) 1.63$\times$±${}_{\pm 0.02}$ 1.11$\times$±${}_{\pm 0.02}$ 2.41$\times$±${}_{\pm 0.04}$ 1.27$\times$±${}_{\pm 0.03}$ 1.70$\times$±${}_{\pm 0.03}$ 1.66$\times$±${}_{\pm 0.04}$ 59.42±${}_{\pm 0.55}$ 1.62$\times$
SpS Leviathan et al. (2023) 1.92$\times$±${}_{\pm 0.04}$ 1.33$\times$±${}_{\pm 0.02}$ 1.93$\times$±${}_{\pm 0.01}$ 1.81$\times$±${}_{\pm 0.04}$ 1.84$\times$±${}_{\pm 0.00}$ 1.76$\times$±${}_{\pm 0.01}$ 64.85±${}_{\pm 0.70}$ 1.77$\times$
Medusa Cai et al. (2024) 1.65$\times$±${}_{\pm 0.03}$ 1.41$\times$±${}_{\pm 0.02}$ 1.33$\times$±${}_{\pm 0.01}$ 1.44$\times$±${}_{\pm 0.03}$ 1.69$\times$±${}_{\pm 0.01}$ 1.29$\times$±${}_{\pm 0.02}$ 54.30±${}_{\pm 0.34}$ 1.48$\times$
EAGLE Li et al. (2024) 2.35$\times$±${}_{\pm 0.03}$ 1.79$\times$±${}_{\pm 0.03}$ 2.04$\times$±${}_{\pm 0.02}$ 1.96$\times$±${}_{\pm 0.03}$ 2.44$\times$±${}_{\pm 0.02}$ 1.80$\times$±${}_{\pm 0.03}$ 76.30±${}_{\pm 0.36}$ 2.08$\times$
$T=1$ Autoregressive Decoding 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 36.24±${}_{\pm 0.43}$ 1.00$\times$
REST He et al. (2023) 1.43$\times$±${}_{\pm 0.01}$ 1.19$\times$±${}_{\pm 0.02}$ 1.24$\times$±${}_{\pm 0.00}$ 1.36$\times$±${}_{\pm 0.02}$ 1.34$\times$±${}_{\pm 0.02}$ 1.61$\times$±${}_{\pm 0.02}$ 49.04±${}_{\pm 0.30}$ 1.35$\times$
SpS Leviathan et al. (2023) 1.55$\times$±${}_{\pm 0.01}$ 1.20$\times$±${}_{\pm 0.01}$ 1.57$\times$±${}_{\pm 0.01}$ 1.54$\times$±${}_{\pm 0.03}$ 1.56$\times$±${}_{\pm 0.03}$ 1.52$\times$±${}_{\pm 0.02}$ 53.94±${}_{\pm 0.43}$ 1.49$\times$
EAGLE Li et al. (2024) 1.79$\times$±${}_{\pm 0.02}$ 1.61$\times$±${}_{\pm 0.03}$ 1.74$\times$±${}_{\pm 0.03}$ 1.66$\times$±${}_{\pm 0.04}$ 1.95$\times$±${}_{\pm 0.06}$ 1.63$\times$±${}_{\pm 0.03}$ 62.88±${}_{\pm 0.54}$ 1.74$\times$
@@
Appendix D Details of Main Experimental Results
The detailed results of our main analysis are shown in Table 5, including the experimental settings of greedy decoding ($T=0$) and nucleus sampling ($T=1$). The findings indicate that EAGLE Li et al. (2024) excels across various Spec-Bench subtasks, achieving an overall speedup ranging from 1.6$\times$ to 2.4$\times$. PLD Saxena (2023) shows notable efficiency in scenarios where the input and output have a significant overlap. For instance, the speedup ratio of PLD increases from 1.27$\times$ in the question answering subtask to 1.66$\times$ in the retrieval-augmented generation subtask, highlighting its effectiveness when the input includes relevant documents. Notably, most methods achieve a suboptimal speedup on the translation subtask. We suspect that it is due to the potential lack of multilingual data in the pretraining corpora.
@@ E.1 Computational Devices
Vicuna-7B Autoregressive Decoding 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 40.24±${}_{\pm 0.30}$ 1.00$\times$
Lookahead Fu et al. (2024) 1.95$\times$±${}_{\pm 0.01}$ 1.61$\times$±${}_{\pm 0.05}$ 1.63$\times$±${}_{\pm 0.02}$ 1.73$\times$±${}_{\pm 0.04}$ 2.16$\times$±${}_{\pm 0.04}$ 1.50$\times$±${}_{\pm 0.00}$ 71.20±${}_{\pm 1.30}$ 1.77$\times$
REST He et al. (2023) 1.72$\times$±${}_{\pm 0.06}$ 1.38$\times$±${}_{\pm 0.05}$ 1.46$\times$±${}_{\pm 0.04}$ 1.80$\times$±${}_{\pm 0.04}$ 1.31$\times$±${}_{\pm 0.03}$ 1.87$\times$±${}_{\pm 0.06}$ 63.81±${}_{\pm 1.00}$ 1.59$\times$
PLD Saxena (2023) 1.67$\times$±${}_{\pm 0.03}$ 1.06$\times$±${}_{\pm 0.03}$ 2.59$\times$±${}_{\pm 0.06}$ 1.16$\times$±${}_{\pm 0.03}$ 1.63$\times$±${}_{\pm 0.03}$ 1.83$\times$±${}_{\pm 0.02}$ 66.61±${}_{\pm 1.15}$ 1.66$\times$
SpS Leviathan et al. (2023) 1.78$\times$±${}_{\pm 0.03}$ 1.19$\times$±${}_{\pm 0.02}$ 1.78$\times$±${}_{\pm 0.03}$ 1.58$\times$±${}_{\pm 0.03}$ 1.54$\times$±${}_{\pm 0.02}$ 1.69$\times$±${}_{\pm 0.02}$ 64.07±${}_{\pm 0.41}$ 1.59$\times$
Medusa Cai et al. (2024) 2.79$\times$±${}_{\pm 0.07}$ 2.36$\times$±${}_{\pm 0.07}$ 2.14$\times$±${}_{\pm 0.04}$ 2.36$\times$±${}_{\pm 0.08}$ 2.77$\times$±${}_{\pm 0.08}$ 2.05$\times$±${}_{\pm 0.01}$ 97.27±${}_{\pm 2.04}$ 2.42$\times$
EAGLE Li et al. (2024) 2.75$\times$±${}_{\pm 0.05}$ 2.08$\times$±${}_{\pm 0.05}$ 2.32$\times$±${}_{\pm 0.05}$ 2.23$\times$±${}_{\pm 0.03}$ 2.79$\times$±${}_{\pm 0.04}$ 2.15$\times$±${}_{\pm 0.01}$ 96.23±${}_{\pm 1.15}$ 2.39$\times$
Vicuna-13B Autoregressive Decoding 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 31.38±${}_{\pm 0.22}$ 1.00$\times$
Lookahead Fu et al. (2024) 1.57$\times$±${}_{\pm 0.01}$ 1.34$\times$±${}_{\pm 0.01}$ 1.39$\times$±${}_{\pm 0.00}$ 1.40$\times$±${}_{\pm 0.01}$ 1.82$\times$±${}_{\pm 0.02}$ 1.32$\times$±${}_{\pm 0.01}$ 46.42±${}_{\pm 0.12}$ 1.48$\times$
REST He et al. (2023) 1.68$\times$±${}_{\pm 0.01}$ 1.31$\times$±${}_{\pm 0.05}$ 1.51$\times$±${}_{\pm 0.01}$ 1.67$\times$±${}_{\pm 0.02}$ 1.29$\times$±${}_{\pm 0.00}$ 1.96$\times$±${}_{\pm 0.01}$ 48.89±${}_{\pm 0.26}$ 1.56$\times$
PLD Saxena (2023) 1.53$\times$±${}_{\pm 0.02}$ 1.08$\times$±${}_{\pm 0.01}$ 2.25$\times$±${}_{\pm 0.00}$ 1.09$\times$±${}_{\pm 0.02}$ 1.65$\times$±${}_{\pm 0.03}$ 1.72$\times$±${}_{\pm 0.00}$ 48.42±${}_{\pm 0.17}$ 1.54$\times$
SpS Leviathan et al. (2023) 1.73$\times$±${}_{\pm 0.02}$ 1.25$\times$±${}_{\pm 0.02}$ 1.76$\times$±${}_{\pm 0.00}$ 1.53$\times$±${}_{\pm 0.01}$ 1.68$\times$±${}_{\pm 0.00}$ 1.73$\times$±${}_{\pm 0.00}$ 50.48±${}_{\pm 0.28}$ 1.61$\times$
Medusa Cai et al. (2024) 2.39$\times$±${}_{\pm 0.02}$ 2.12$\times$±${}_{\pm 0.02}$ 1.92$\times$±${}_{\pm 0.00}$ 2.07$\times$±${}_{\pm 0.02}$ 2.49$\times$±${}_{\pm 0.02}$ 1.88$\times$±${}_{\pm 0.00}$ 67.64±${}_{\pm 0.07}$ 2.16$\times$
EAGLE Li et al. (2024) 2.88$\times$±${}_{\pm 0.05}$ 2.24$\times$±${}_{\pm 0.04}$ 2.52$\times$±${}_{\pm 0.03}$ 2.24$\times$±${}_{\pm 0.04}$ 2.90$\times$±${}_{\pm 0.03}$ 2.34$\times$±${}_{\pm 0.01}$ 79.35±${}_{\pm 1.18}$ 2.53$\times$
Vicuna-33B Autoregressive Decoding 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 1.00$\times$±${}_{\pm 0.00}$ 16.34±${}_{\pm 0.01}$ 1.00$\times$
Lookahead Fu et al. (2024) 1.46$\times$±${}_{\pm 0.00}$ 1.21$\times$±${}_{\pm 0.00}$ 1.32$\times$±${}_{\pm 0.00}$ 1.29$\times$±${}_{\pm 0.00}$ 1.71$\times$±${}_{\pm 0.00}$ 1.28$\times$±${}_{\pm 0.00}$ 22.58±${}_{\pm 0.08}$ 1.38$\times$
REST He et al. (2023) 1.71$\times$±${}_{\pm 0.01}$ 1.39$\times$±${}_{\pm 0.00}$ 1.57$\times$±${}_{\pm 0.00}$ 1.69$\times$±${}_{\pm 0.01}$ 1.34$\times$±${}_{\pm 0.01}$ 1.89$\times$±${}_{\pm 0.00}$ 25.98±${}_{\pm 0.07}$ 1.59$\times$
PLD Saxena (2023) 1.45$\times$±${}_{\pm 0.00}$ 1.06$\times$±${}_{\pm 0.00}$ 1.98$\times$±${}_{\pm 0.00}$ 1.07$\times$±${}_{\pm 0.00}$ 1.54$\times$±${}_{\pm 0.00}$ 1.43$\times$±${}_{\pm 0.00}$ 23.07±${}_{\pm 0.01}$ 1.41$\times$
SpS Leviathan et al. (2023) 1.79$\times$±${}_{\pm 0.00}$ 1.31$\times$±${}_{\pm 0.00}$ 1.80$\times$±${}_{\pm 0.00}$ 1.57$\times$±${}_{\pm 0.00}$ 1.73$\times$±${}_{\pm 0.00}$ 1.69$\times$±${}_{\pm 0.00}$ 26.89±${}_{\pm 0.03}$ 1.65$\times$
Medusa Cai et al. (2024) 2.22$\times$±${}_{\pm 0.00}$ 1.95$\times$±${}_{\pm 0.00}$ 1.85$\times$±${}_{\pm 0.00}$ 1.87$\times$±${}_{\pm 0.01}$ 2.32$\times$±${}_{\pm 0.01}$ 1.84$\times$±${}_{\pm 0.00}$ 32.92±${}_{\pm 0.06}$ 2.01$\times$
EAGLE Li et al. (2024) 2.81$\times$±${}_{\pm 0.00}$ 2.14$\times$±${}_{\pm 0.00}$ 2.53$\times$±${}_{\pm 0.00}$ 2.19$\times$±${}_{\pm 0.00}$ 3.01$\times$±${}_{\pm 0.00}$ 2.31$\times$±${}_{\pm 0.00}$ 40.91±${}_{\pm 0.03}$ 2.50$\times$

### v3

@@ B.1 Details of Spec-Bench
Moreover, Spec-Bench incorporates three further subtasks – translation, question answering, and mathematical reasoning – to provide a thorough evaluation of Speculative Decoding’s speedup capabilities in diverse contexts. We utilized WMT14 DE-EN, Natural Questions, and GSM8K Cobbe et al. (2021) as the primary datasets for these tasks, respectively. We randomly selected 80 instances from each subtask’s test set for evaluation. The detailed composition is summarized in Table 4.
@@
$T=0$ Autoregressive Decoding 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 36.74±0.31 1.00$\times$
Lookahead Fu et al. (2024) 1.15$\times$±0.01 0.98$\times$±0.02 1.07$\times$±0.02 1.06$\times$±0.02 1.32$\times$±0.02 1.03$\times$±0.02 40.64±0.26 1.11$\times$
REST He et al. (2023) 1.49$\times$±0.02 1.23$\times$±0.04 1.26$\times$±0.03 1.39$\times$±0.04 1.34$\times$±0.03 1.71$\times$±0.05 51.12±0.78 1.39$\times$
PLD Saxena (2023) 1.63$\times$±0.02 1.11$\times$±0.02 2.41$\times$±0.04 1.27$\times$±0.03 1.70$\times$±0.03 1.66$\times$±0.04 59.42±0.55 1.62$\times$
SpS Leviathan et al. (2023) 1.92$\times$±0.04 1.33$\times$±0.02 1.93$\times$±0.01 1.81$\times$±0.04 1.84$\times$±0.00 1.76$\times$±0.01 64.85±0.70 1.77$\times$
Medusa Cai et al. (2024) 1.65$\times$±0.03 1.41$\times$±0.02 1.33$\times$±0.01 1.44$\times$±0.03 1.69$\times$±0.01 1.29$\times$±0.02 54.30±0.34 1.48$\times$
EAGLE Li et al. (2024) 2.35$\times$±0.03 1.79$\times$±0.03 2.04$\times$±0.02 1.96$\times$±0.03 2.44$\times$±0.02 1.80$\times$±0.03 76.30±0.36 2.08$\times$
$T=1$ Autoregressive Decoding 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 36.24±0.43 1.00$\times$
REST He et al. (2023) 1.43$\times$±0.01 1.19$\times$±0.02 1.24$\times$±0.00 1.36$\times$±0.02 1.34$\times$±0.02 1.61$\times$±0.02 49.04±0.30 1.35$\times$
SpS Leviathan et al. (2023) 1.55$\times$±0.01 1.20$\times$±0.01 1.57$\times$±0.01 1.54$\times$±0.03 1.56$\times$±0.03 1.52$\times$±0.02 53.94±0.43 1.49$\times$
EAGLE Li et al. (2024) 1.79$\times$±0.02 1.61$\times$±0.03 1.74$\times$±0.03 1.66$\times$±0.04 1.95$\times$±0.06 1.63$\times$±0.03 62.88±0.54 1.74$\times$
@@
B.2 Implementation Details
Appendix C Details of Main Experimental Results
The detailed results of our main analysis are shown in Table 5, including the experimental settings of greedy decoding ($T=0$) and speculative sampling ($T=1$). The findings indicate that EAGLE Li et al. (2024) excels across various Spec-Bench subtasks, achieving an overall speedup ranging from 1.6$\times$ to 2.4$\times$. PLD Saxena (2023) shows notable efficiency in scenarios where the input and output have a significant overlap. For instance, the speedup ratio of PLD increases from 1.27$\times$ in the question answering subtask to 1.66$\times$ in the retrieval-augmented generation subtask, highlighting its effectiveness when the input includes relevant documents. Notably, most methods achieve a suboptimal speedup on the translation subtask. We suspect that it is due to the potential lack of multilingual data in the pretraining corpora.
@@ Appendix C Details of Main Experimental Results
Vicuna-7B Autoregressive Decoding 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 40.24±0.30 1.00$\times$
Lookahead Fu et al. (2024) 1.95$\times$±0.01 1.61$\times$±0.05 1.63$\times$±0.02 1.73$\times$±0.04 2.16$\times$±0.04 1.50$\times$±0.00 71.20±1.30 1.77$\times$
REST He et al. (2023) 1.72$\times$±0.06 1.38$\times$±0.05 1.46$\times$±0.04 1.80$\times$±0.04 1.31$\times$±0.03 1.87$\times$±0.06 63.81±1.00 1.59$\times$
PLD Saxena (2023) 1.67$\times$±0.03 1.06$\times$±0.03 2.59$\times$±0.06 1.16$\times$±0.03 1.63$\times$±0.03 1.83$\times$±0.02 66.61±1.15 1.66$\times$
SpS Leviathan et al. (2023) 1.78$\times$±0.03 1.19$\times$±0.02 1.78$\times$±0.03 1.58$\times$±0.03 1.54$\times$±0.02 1.69$\times$±0.02 64.07±0.41 1.59$\times$
Medusa Cai et al. (2024) 2.79$\times$±0.07 2.36$\times$±0.07 2.14$\times$±0.04 2.36$\times$±0.08 2.77$\times$±0.08 2.05$\times$±0.01 97.27±2.04 2.42$\times$
EAGLE Li et al. (2024) 2.75$\times$±0.05 2.08$\times$±0.05 2.32$\times$±0.05 2.23$\times$±0.03 2.79$\times$±0.04 2.15$\times$±0.01 96.23±1.15 2.39$\times$
Vicuna-13B Autoregressive Decoding 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 31.38±0.22 1.00$\times$
Lookahead Fu et al. (2024) 1.57$\times$±0.01 1.34$\times$±0.01 1.39$\times$±0.00 1.40$\times$±0.01 1.82$\times$±0.02 1.32$\times$±0.01 46.42±0.12 1.48$\times$
REST He et al. (2023) 1.68$\times$±0.01 1.31$\times$±0.05 1.51$\times$±0.01 1.67$\times$±0.02 1.29$\times$±0.00 1.96$\times$±0.01 48.89±0.26 1.56$\times$
PLD Saxena (2023) 1.53$\times$±0.02 1.08$\times$±0.01 2.25$\times$±0.00 1.09$\times$±0.02 1.65$\times$±0.03 1.72$\times$±0.00 48.42±0.17 1.54$\times$
SpS Leviathan et al. (2023) 1.73$\times$±0.02 1.25$\times$±0.02 1.76$\times$±0.00 1.53$\times$±0.01 1.68$\times$±0.00 1.73$\times$±0.00 50.48±0.28 1.61$\times$
Medusa Cai et al. (2024) 2.39$\times$±0.02 2.12$\times$±0.02 1.92$\times$±0.00 2.07$\times$±0.02 2.49$\times$±0.02 1.88$\times$±0.00 67.64±0.07 2.16$\times$
EAGLE Li et al. (2024) 2.88$\times$±0.05 2.24$\times$±0.04 2.52$\times$±0.03 2.24$\times$±0.04 2.90$\times$±0.03 2.34$\times$±0.01 79.35±1.18 2.53$\times$
Vicuna-33B Autoregressive Decoding 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 1.00$\times$±0.00 16.34±0.01 1.00$\times$
Lookahead Fu et al. (2024) 1.46$\times$±0.00 1.21$\times$±0.00 1.32$\times$±0.00 1.29$\times$±0.00 1.71$\times$±0.00 1.28$\times$±0.00 22.58±0.08 1.38$\times$
REST He et al. (2023) 1.71$\times$±0.01 1.39$\times$±0.00 1.57$\times$±0.00 1.69$\times$±0.01 1.34$\times$±0.01 1.89$\times$±0.00 25.98±0.07 1.59$\times$
PLD Saxena (2023) 1.45$\times$±0.00 1.06$\times$±0.00 1.98$\times$±0.00 1.07$\times$±0.00 1.54$\times$±0.00 1.43$\times$±0.00 23.07±0.01 1.41$\times$
SpS Leviathan et al. (2023) 1.79$\times$±0.00 1.31$\times$±0.00 1.80$\times$±0.00 1.57$\times$±0.00 1.73$\times$±0.00 1.69$\times$±0.00 26.89±0.03 1.65$\times$
Medusa Cai et al. (2024) 2.22$\times$±0.00 1.95$\times$±0.00 1.85$\times$±0.00 1.87$\times$±0.01 2.32$\times$±0.01 1.84$\times$±0.00 32.92±0.06 2.01$\times$
EAGLE Li et al. (2024) 2.81$\times$±0.00 2.14$\times$±0.00 2.53$\times$±0.00 2.19$\times$±0.00 3.01$\times$±0.00 2.31$\times$±0.00 40.91±0.03 2.50$\times$